OpenStudy (anonymous):

The length and width of a rectangle are 17 cm and 9 cm. If the dimensions are tripled, by exactly how much has the perimeter of the rectangle increased? (Hint: Use the Numeric Method) 26 cm 104 cm 78 cm 52 cm

OpenStudy (mertsj):

104 cm

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Normal Perimeter = 2(l + b) cm = 2(17+9) cm = 52 cm If tripled perimeters= 2(51 + 27) cm = 2(78) cm = 156 cm So therefore, 156 cm - 52 cm = 104 cm.